# Harnessing natural and mechanical airflows for surface-based atmospheric pollutant removal Centre for Climate Repair, Cambridge, UK

### v4

@@
Harnessing natural and mechanical airflows for surface-based atmospheric pollutant removal
Centre for Climate Repair, Cambridge, UK
Department of Mechanical Engineering, Stanford University, Stanford, USA
May 22, 2025)
@@
Removal strategies for atmospheric pollutants are increasingly being considered to mitigate global warming and improve public health. However, surface-based removal techniques, such as sorption, catalysis and filtration, are often limited by pollutant transport and removal rate constraints. We evaluate the atmospheric pollutant transport to surfaces and assess the potential of surface-based removal technologies for applications in airflow through cities, HVAC systems and over vehicles. If these removal technologies are applied to their surfaces, cities, solar farms, HVAC systems and filters can achieve atmospheric pollutant removal rates that exceed 1 GtCO2e annually (20-year GWP). Cities have the highest atmospheric pollutant removal potential, with estimates averaging 30 GtCO2, 0.06 GtCH4, 0.0001 GtPM2.5, 0.007 GtNO annually. HVAC filters can achieve atmospheric pollutant removal costs as low as $300 per tCO2e removed when sorption or catalyst technologies are incorporated into their fibre sheets, outperforming the $2000 per tCO2e removal cost when these technologies are applied to city surfaces. This estimate is based on the literature values for these technologies’ costs per square meter. However, our calculations indicate that optimising catalyst properties and surface coverage could lower the cost estimates to below $100 per tCO2e across these applications. These findings demonstrate that integrating surface-based pollutant removal technologies into infrastructure may offer a scalable pathway to advance climate and health objectives.
@@
Greenhouse gases (GHGs) such as carbon dioxide (CO2), methane (CH4) and nitrous oxide (N2O) continue to rise at significant rates, presenting sustainability and health challenges by accelerating climate change [22, 21]. In 2023, CO2 levels surpassed 420 ppm, driven by fossil fuel use and deforestation, contributing to biodiversity loss and environmental degradation [35, 22]. The situation is compounded by CH4 and N2O, which, although emitted in smaller quantities, have global warming potentials (GWPs) 84 and 273 times that of CO2 over a 20-year timescale. These potent GHGs are emitted by anthropogenic sources such as agriculture and fossil fuel processing [3, 56, 12, 38]. However, unlike CO2, CH4 and N2O have biogenic sources that are difficult to control without disrupting agriculture and ecosystems. In addition to GHGs, atmospheric pollutants such as particulate matter (PM), sulphur dioxide (SO2) and nitrogen oxides (NO) pose environmental and medical risks [27, 30]. PM, particularly particles smaller than 2.5 m (PM2.5), is a leading cause of respiratory and cardiovascular diseases, causing millions of premature deaths annually [47, 30, 33]. PM is emitted from anthropogenic sources, such as fossil fuel processing, and natural sources, like wildfires; SO2 and NO have similar sources, but contribute to acid rain and ground-level ozone [7, 2, 5].
@@
Given the diffuse nature of anthropogenic, biogenic and natural sources, coupled with the insufficient progress in emissions reductions, there is growing interest in atmospheric pollutant removal [35, 22, 12, 38, 30, 33, 7, 2, 5]. Atmospheric pollutant removal can generally be divided into two categories: surface-based approaches [39, 25, 26, 16, 23, 31, 57, 55], which rely on the interaction between pollutants and surfaces, and airborne methods, which use radicals or surface chemistry to break down pollutants in the air [54, 36]. While airborne methods show promise, this study focuses on surface-based removal methods, given that these approaches are more widely explored and developed. Despite the interest in surface-based atmospheric pollutant removal [39, 25, 26, 16, 23, 31, 57, 55], the scale of potential removal remains unexplored. This study aims to assess the potential of various surface-based strategies by quantifying the scales at which pollutants can be removed from the atmosphere. Nonetheless, the methods developed here apply to other pollutants, offering a framework for quantifying the removal potential across contexts.
@@
Direct air capture (DAC) technologies, often sorbent (e.g., amines) or solvent-based (e.g., potassium hydroxide), can capture CO2 through chemical adsorption and absorption processes [39, 8]. For DAC, the fraction of pollutant molecules removed upon contact with the surface, known as the removal efficiency, ranges from to [46]. However, this process requires substantial energy and involves high operational costs [29, 28]. Membrane-based systems, which separate CO2 through selectively permeable barriers, offer greater energy efficiency than DAC, but face challenges related to initial costs and scalability [25]. Hence, while both methods show promise, their widespread application has been limited. In contrast, leveraging pre-existing natural and mechanical airflows, where the associated energy demands are already accounted for, may enable scalable and cost-effective surface-based methods for atmospheric pollutant removal.
@@
Catalytic oxidation can convert CH4 into CO2 and water (H2O), reducing its greenhouse impact; however, performance optimisation across environments remains challenging [26]. Reported CH4 removal efficiencies for thermal, photo and electrocatalysts vary widely, from to , reflecting differences across catalysts and experimental conditions [50, 32]. Recent cost analyses suggest that using photocatalytic oxidation for atmospheric CH4 removal is prohibitively expensive due to the high flow rates required for effective removal [36, 32, 18]. Alternatively, atmospheric radicals can break down CH4 but at slow, condition-dependent rates [54, 36] and methanotrophic bacteria can remove CH4, albeit under controlled conditions [16]. This study examines catalytic oxidation as a strategy for atmospheric pollutant removal, exploring applications with varying surface fluxes and areas for scalable implementation.
Similar to CH4, N2O can be removed via catalytic reduction, which converts N2O into nitrogen (N2) and oxygen (O2) [36]. Another surface-based removal process is biological denitrification, where bacteria reduce N2O to N2 under anaerobic conditions [49].
@@
High-efficiency particulate air (HEPA) filters are effective at capturing PM [23], alongside other methods such as electrostatic precipitators and adsorption materials such as activated carbon [31]. These technologies offer removal efficiencies of to [42], but require periodic regeneration or replacement due to saturation [24]. PM is also removed from the atmosphere through dry (e.g., gravitational settling, surface absorption) and wet deposition (e.g., scavenging by rain), with removal rates varying by region and environmental conditions [45, 53, 15]. While these mechanisms for PM removal are well-established, our study will estimate how much PM2.5 is removed from the atmosphere by filtration technologies and dry deposition via surface absorption, comparing these with the potential removal rates of other pollutants.
Similar to CH4 and N2O, SO2 and NO can be removed using catalysis, which converts these pollutants into sulfuric acid (H2SO4) and N2 [57]. Another removal method is scrubbing, where SO2 and NO are removed from gas streams by passing them through chemically reactive liquid solutions [55].
@@
In this study, we examine approaches that could utilise the substantial volumes of air transported in urban areas, heating, ventilation and air conditioning (HVAC) systems, and transportation networks, where pollutants interact with surfaces to enable their removal. Rather than focusing on the performance of specific surface-based technologies that remove specific pollutants, this study generally evaluates the scalability of atmospheric pollutant removal via surface-based interactions through transport limitations. We use a combination of theoretical models, empirical relations, global measurements and industry standards to estimate the atmospheric pollutant flow rate to the surfaces of various natural and mechanical systems. By leveraging removal efficiencies measured in the literature [46, 50, 32, 42], we assess the projected removal rates and technology costs for sorption, catalysis and filtration, for a range of environmental and flow conditions. We compare the potential atmospheric pollutant removal rates with intergovernmental targets and existing removal mechanisms, evaluating the scalability and cost-effectiveness of these surface-based technologies.
@@ Results
Natural (e.g., cities, solar farms), internal (e.g., HVAC, combustion, DAC systems) and external environments (e.g., aeroplanes, trains, automobiles) offer significant opportunities for atmospheric pollutant removal. As shown in Fig. 1a and detailed in Supplementary Sec. 3, these environments consist of substantial airflow volumes that drive atmospheric pollutant fluxes to surfaces. Average streamwise flow rates are approximately m3/s through natural environments, m3/s in internal systems and m3/s around external systems. Theoretical predictions, empirical data, global measurements and industry standards provide a range of estimates in Fig. 1a, spanning from to m3/s for cities, to m3/s for HVAC systems and to m3/s for automobiles. Key parameters, including length, surface area, velocity and the number of environments are summarised in Fig. 1b–g, Supplementary Tabs. 2–4 and Secs. 2.1–2.3. The parameter distributions not only capture uncertainty but reveal that surface area and airflow velocity are the dominant factors influencing pollutant removal efficiency across applications.
@@
Fig. 2a shows normalised atmospheric pollutant fluxes to the surfaces of natural environments (columns 1–2) in cities and solar farms. GHG fluxes exceed other pollutant fluxes due to higher diffusivities and atmospheric concentrations (Fig. 2c). Atmospheric pollutant fluxes to surfaces are also influenced by length scales, with cities exhibiting average surface fluxes of mol/m2s for CO2, while solar farms achieve average surface fluxes of mol/m2s for CO2. Smaller length scales reduce the boundary layer thickness, which increases the turbulence intensity and improves the atmospheric pollutant transfer to surfaces. As a result, in these natural environments, atmospheric pollutant fluxes to surfaces are averaged across scales. Using empirical formulae detailed in the Methods, average atmospheric pollutant fluxes at the repeating unit scale (e.g., to the surface of individual buildings, m-1) are larger than those at the combined unit scale (e.g., to the surface of entire cities, m-1) due to steeper concentration gradients (Figs. 2a, d, e).
@@
Figs. 2a–b highlight the enhanced atmospheric pollutant transport to the surfaces of internal airflows (columns 3–5) in HVAC, combustion, DAC systems (Fig. 2a) and HVAC filters (Fig. 2b). Comparing HVAC, combustion and DAC systems, which show average surface fluxes of , and mol/m2s for CO2 in Fig. 2c, demonstrates that higher airflow velocities enhance atmospheric pollutant transport. The estimated value for DAC is comparable to mol/m2s estimated in McQueen et al. [28]. At the duct wall, the rate at which pollutants move perpendicular to the surface (normal flux) depends on diffusion and is proportional to (e.g., an average of mol/m2s for CO2, Fig. 2c, f), where is the diffusivity of the pollutant, is its concentration and is the height of the channel. In contrast, for HVAC filters, the pollutant flux through the porous fibre sheet is proportional to (e.g., an average of mol/m2s for CO2, Fig. 2c, g), where is the airflow velocity through the fibre sheet. Since and are typically small relative to and , is generally larger than , highlighting the benefits of utilising advective over diffusive transport to improve atmospheric pollutant fluxes to surfaces.
@@
Lastly, we evaluate the normalised atmospheric pollutant fluxes to the surfaces of external airflows over aeroplanes, trains and automobiles (columns 6–8, Fig. 2a). Automobiles demonstrate higher pollutant fluxes to surfaces than trains and aeroplanes (e.g., an average of , , mol/m2s for CO2, Fig. 2c), due to smaller characteristic lengths. However, aeroplanes generate more turbulent transport due to their higher velocities, resulting in a pollutant flux that is approximately 25% of the automobile flux across the entire surface.
@@
Fig. 3 quantifies the flow rates of atmospheric CO2, CH4, PM2.5 and NO to all global surfaces. These flow rates are evaluated by multiplying the fluxes from Fig. 2 by the relevant surface area for each application ( in Fig. 2d–g) and the number of each application. Natural and internal airflows (columns 1, 3, 4) with extensive surface areas yield the highest atmospheric pollutant flow rates to their surfaces, e.g., an average of 30 GtCO2/y for cities, 10 GtCO2/y for HVAC systems and 4 GtCO2/y for HVAC filters. Conversely, external airflows and small-surface-area internal systems (columns 5–8) show lower atmospheric pollutant flow rates to their surface, e.g., an average of 0.1 GtCO2/y for aeroplanes, 0.7 GtCO2/y for trains, GtCO2/y for combustion systems and GtCO2/y for DAC systems, despite external airflows exhibiting higher atmospheric pollutant fluxes to their surfaces in Fig. 2a–c.
Integrating surface-based pollutant removal technologies into global infrastructure is likely more readily achieved by incorporating technologies into newly-produced surfaces. Comparing atmospheric pollutant flow rates to total existing surfaces (Fig. 3) with annually produced surfaces (Supplementary Fig. 2), a decrease is observed across applications, with the reduction varying between applications. For example, solar farms and cities have average flow rates of 30 GtCO2/y and 0.08 GtCO2/y to their total surfaces. However, when accounting for annually produced surfaces, the average flow rate to new surfaces in cities (0.2 GtCO2/y) is more significantly reduced than for solar farms (0.03 GtCO2/y). This is largely driven by the global growth of solar farms, expanding at approximately 20% annually, compared with the 4% annual growth of cities.

### v5

# Evaluating the scale of natural and mechanical airflows for surface-based atmospheric pollutant removal

@@
Abstract
@@
Removal strategies for atmospheric pollutants are increasingly being considered to mitigate global warming and improve public health. However, the global potential of surface-based removal techniques (e.g. sorption, catalysis and filtration) has not yet been quantified based on limits of pollutant transport and removal rates. We evaluate the atmospheric pollutant transport to surfaces and assess the potential of surface-based removal technologies for global-scale deployment across a variety of configurations, including air interaction with the built environment, mechanical ventilation and convection systems, and over the global transportation fleet. Cities provide the highest removal potential, with median annual estimates of 30 GtCO2, 0.06 GtCH4, 0.007 GtNO and 0.0001 GtPM2.5 transported to their total surface area. Cities, solar farms, HVAC systems and filters have the potential to exceed 1 GtCO2/yr (1 GtCO2e/yr for CH4, 20-year GWP) of removal when literature-based CO2-sorbent (CH4-catalyst) efficiencies are applied across their total surface area. These values represent theoretical upper bounds and are intended for comparison across applications rather than application-specific deployment. HVAC filters have the potential to achieve materials costs as low as $600 per tCO2 removed ($2000 per tCO2e) when CO2-sorption (CH4-catalyst) technologies are incorporated into their fibre sheets and maintained through routine filter replacement, compared with $3000 per tCO2 ($10000 per tCO2e) for city surfaces, based on the literature values for these technologies’ materials costs. These findings demonstrate that integrating surface-based pollutant removal technologies into infrastructure may offer a pathway to advance climate and health objectives, though further studies are needed to assess their feasibility in application, and application-implementation rates and cost.
keywords:
[…]
Keywords: pollutant removal; cities; HVAC systems; sorption; catalysis; filtration.
Synopsis: Surface-based removal technologies applied to existing infrastructure can enable substantial atmospheric pollutant removal, supporting climate and air quality goals.
1 Introduction
@@
Greenhouse gases (GHGs) such as carbon dioxide (CO2), methane (CH4) and nitrous oxide (N2O) continue to rise at substantial rates, presenting sustainability and health challenges by accelerating climate change 25, 22. In 2023, CO2 levels surpassed 420 ppm, driven by fossil fuel use and deforestation, contributing to biodiversity loss and environmental degradation 41, 25. Although CH4 and N2O are emitted in smaller quantities than CO2, their 20-year global warming potentials (GWPs) are 84 and 273 times greater than CO2, highlighting their disproportionate climate impacts despite lower atmospheric concentrations 25, 22. CO2 is emitted from anthropogenic activities (e.g., fossil fuel combustion, land-use change) and natural processes (e.g., respiration, decomposition), and CH4 and N2O arise from a combination of anthropogenic sources, such as agriculture and fossil fuel processing, and natural microbial processes 4, 62, 12, 43. Many GHG sources are spatially distributed and occur at low concentrations, making them difficult to remove.
In addition to GHGs, atmospheric pollutants such as particulate matter (PM), sulphur dioxide (SO2), nitrogen oxides (NO) and ozone (O3) contribute to a range of environmental and health risks 31, 35. PM, particularly particles smaller than 2.5 m (PM2.5), is a leading cause of respiratory and cardiovascular diseases, responsible for millions of premature deaths each year 53, 35, 39. PM is emitted from anthropogenic sources, such as fossil fuel combustion and processing, and natural sources, like wildfires. It can also form secondarily from precursor gases including SO2 and NO, which have similar sources to primary PM and contribute to acid rain and ground-level O3 7, 3, 5. O3, a secondary pollutant formed photochemically from NO and volatile organic compounds (VOCs), exacerbates respiratory illness and contributes indirectly to radiative forcing, though its atmospheric lifetime is shorter than those of long-lived GHGs 63.
@@
Given the diffuse nature of anthropogenic, biogenic and natural sources, coupled with the insufficient progress in emissions reductions, there is growing interest in atmospheric pollutant removal 41, 25, 12, 43, 35, 39, 7, 3, 5. Atmospheric pollutant removal strategies can broadly be grouped into (i) surface-based approaches 45, 28, 29, 16, 26, 36, 64, 61, which capture or transform pollutants at solid interfaces, and (ii) airborne chemical methods 60, 42, such as radical generation or advanced photochemistry. While airborne methods show promise, this study focuses on surface-based removal methods such as sorption, catalysis and filtration, given that these approaches are more widely explored and developed. Despite the interest in surface-based atmospheric pollutant removal 45, 28, 29, 16, 26, 36, 64, 61, the scale of potential removal remains unexplored. This study quantifies the transport of CO2, CH4, NO and PM2.5 to natural and engineered surfaces to estimate upper-bound removal potentials, providing a comparative framework across environments rather than prescribing deployment pathways. Nonetheless, the methods developed here apply to other pollutants, offering a framework for quantifying the removal potential across contexts.
@@
Sorption technologies, often sorbent (e.g., amines) or solvent-based (e.g., potassium hydroxide), can capture CO2 through chemical adsorption and absorption processes, as applied in direct air capture (DAC) 45, 8. For DAC, the fraction of pollutant molecules removed upon contact with the surface, known as the removal efficiency, ranges from to 52. However, this process requires substantial energy and involves high operational costs 33, 32. Membrane-based systems, which separate CO2 through selectively permeable barriers, offer greater energy efficiency than DAC, but face challenges related to initial costs and scalability 28. Hence, while both methods show promise, their widespread application has been limited. In contrast, leveraging pre-existing natural and mechanical airflows, where the associated energy demands are already accounted for, may enable scalable and cost-effective surface-based methods for atmospheric pollutant removal.
@@
Catalytic oxidation can convert CH4 into CO2 and water (H2O), reducing its greenhouse impact; however, performance optimisation across environments remains challenging 29. Reported CH4 removal efficiencies for thermal, photo and electrocatalysts vary widely, from to , reflecting differences across catalysts and experimental conditions 56, 37. Recent cost analyses suggest that using photocatalytic oxidation for atmospheric CH4 removal is prohibitively expensive due to the high flow rates required for effective removal 42, 37, 18. Further analysis shows that targeting higher-concentration CH4 sources can yield meaningful climate ( ppm) and cost benefits ( ppm), with limited feasibility at ambient CH4 levels due to efficiency and energy constraints 1. Alternatively, atmospheric radicals can break down CH4 but at slow, condition-dependent rates 60, 42 and methanotrophic bacteria can remove CH4, albeit under controlled conditions 16. This study examines catalytic oxidation as a strategy for atmospheric pollutant removal, exploring applications with varying surface fluxes and areas for scalable implementation.
@@
High-efficiency particulate air (HEPA) filters are widely used for PM capture, though their performance depends on particle size distributions, pressure drop constraints and replacement cycles 26; additional coatings can extend their removal potential to some gaseous pollutants. Other methods for removing PM include electrostatic precipitators and adsorption materials such as activated carbon 36. These technologies offer removal efficiencies of to 48, but require periodic regeneration or replacement due to saturation 27. PM is also removed from the atmosphere through dry (e.g., gravitational settling, surface absorption) and wet deposition (e.g., scavenging by rain), with removal rates varying by region and environmental conditions 51, 59, 15. While these mechanisms for PM removal are well-established, our study will estimate how much PM2.5 is removed from the atmosphere by filtration technologies and dry deposition via surface absorption, comparing these with the potential removal rates of other pollutants.
Similar to CH4, N2O can be removed via catalytic reduction, which converts N2O into nitrogen (N2) and oxygen (O2) 42. Another surface-based removal process is biological denitrification, where bacteria reduce N2O to N2 under anaerobic conditions 55. SO2, NO and O3 can be transformed through catalytic processes (e.g., selective catalytic reduction for NO, oxidation pathways for SO2, photocatalysis for O3), with variability in efficiency, by-products and technological maturity 64. Another removal method is scrubbing, where SO2 and NO are removed from gas streams by passing them through chemically reactive liquid solutions 61. Note that simultaneous removal of distinct pollutants would require multiple technologies operating in parallel or complementary configurations, as no single removal technology can achieve all pollutant removal pathways simultaneously.
@@
In this study, we examine approaches that could utilise the substantial volumes of air transported in urban areas, heating, ventilation and air conditioning (HVAC) systems, and transportation networks, where pollutants interact with surfaces to enable their removal. Rather than focusing on the performance of specific surface-based technologies that remove specific pollutants, this study generally evaluates the scalability of atmospheric pollutant removal via surface-based interactions through transport limitations. We use a combination of theoretical models, empirical relations, global measurements and industry standards to estimate the atmospheric pollutant flow rate to the surfaces of various natural and mechanical systems. By leveraging removal efficiencies measured in the literature 52, 56, we assess the projected removal rates and materials costs for sorption, catalysis and filtration, for a range of environmental and flow conditions. We compare the potential atmospheric pollutant removal rates with intergovernmental targets and existing removal mechanisms, evaluating the scalability and cost-effectiveness of these surface-based technologies.
2 Methods
The methods outlined in this section are used to calculate the pollutant flux or flow rate to the surface of different natural and mechanical environments. These fluxes or flow rates are independent of the removal technology and indicate the rate at which a given atmospheric pollutant could be delivered to the surface of a given environment (to potentially be removed). These methods are divided into scaling theories; empirical relationships; energy, power and drag measurements; industry standards; and fully-developed profiles.
2.1 Scaling theory
To characterise atmospheric pollutant transport dynamics across various systems, we define velocity (, m/s), length (, m) and surface area (, m2) scales specific to each flow application (e.g., cities, HVAC systems, transport). We also estimate the number of occurrences () for each flow type to ensure realistic scaling for global estimations. The limitations of this scaling approach and additional context are provided in Section S3.
The thickness of the turbulent () boundary layer (, m) is evaluated using , where the Reynolds number 47 and is the kinematic viscosity (m2/s). The normal pollutant flux (mol/m2/s) at the surface is determined by the diffusive flux, , as the normal velocity at the surface is zero. The normal pollutant flux is simplified here using Fick’s law, which assumes diffusion driven solely by concentration gradients without accounting for molecular interactions or multicomponent effects. This normal pollutant flux depends on the diffusivity (m2/s), concentration (mol/m3) and the boundary layer thickness (m) 13, where the Schmidt number . An effective diffusivity (m2/s) is included to account for enhanced mixing, which alters pollutant transport rates 40. We will assume well-mixed concentrations and constant effective diffusivities throughout. Their spatio-temporal variations due to environmental factors and sources will be pursued in future work.
The formulas in (1)–(2) provide a means to estimate atmospheric pollutant transport through, and to the surface of, global environments. As shown in Fig. S1, we also calculate the global streamwise flow rate through the boundary layers of the world’s cities (solar farms) using ERA5, SimpleMaps and Dunnett et al. 17, 50, 11. For some applications (e.g., cities, solar farms), multiple characteristic length scales are applied (e.g., buildings, solar panels), whereas for others (e.g., ducts, aeroplanes), only a single length scale is used. For applications with a single length scale, (1)–(2) are evaluated using parameter distributions from Tables S1–S3. For multiple length scale applications, (1)–(2) are evaluated separately for each length scale and the results are combined into a single distribution.
2.2 Empirical relationships
Empirical models can be used to link flow characteristics with the potential for pollutant removal 13. These relationships have been applied in CH4 and N2O removal studies via photocatalysis 56, 42. These relationships are evaluated using , and , supplemented with measured values such as wall shear stresses ( in kg/ms2), mass transfer coefficients ( in m/s) and heat transfer coefficients ( in W/m2K). A detailed discussion of these empirical relationships and their applications can be found in Section S3.
Analogies exist between key dimensionless numbers: the Nusselt number (), the Sherwood number () and the drag coefficient (), where is the thermal conductivity (W/mK) and is the fluid density (kg/m3) 13. These analogies allow for the formulation of the normal pollutant flow rate to the surface of each application as follows
Following Tsopelakou et al. 56, we can use (3) to estimate the streamwise flow rate through each application as
where is the wetted perimeter and is the cross-sectional area of the channel or boundary layer. Eqs. (3)–(4) are evaluated using the parameter distributions listed in Tabs. S1–S3, and the resulting distribution is then combined with the other methods from this section.
2.3 Energy, power and drag measurements
We now outline formulas to determine the flow rate based on energy, power and drag measurements. More details are given in Section S3.
For external flows over urban areas or solar farms, we establish a control volume with a length and height . In the case of uni-directional, steady-state flow over a surface with a heat flux (W/m2), energy conservation dictates that the energy entering the control volume must balance with the energy exiting it 13. This relationship can be expressed as
where (m2/s2K) is the specific heat capacity of air at constant pressure and (K) is the temperature difference between the surface and the surrounding atmosphere. The flow rates are related using the scaling theory discussed in Section S3, such that .
where is the leakage factor 46.
For external airflows acting on transportation systems, the streamwise flow rate through the boundary layer can be estimated using the drag force acting on the aeroplane, train or automobile 58. This is expressed by the equation
where (N) is the drag force. Eqs. (5)–(7) are evaluated using the parameter distributions listed in Tabs. S1–S3, and the resulting distribution is then combined with the other methods from this section.
2.4 Industry standards
For internal airflows through HVAC ducting, the streamwise flow rate can be evaluated using three estimates 2. First, can be expressed in terms of the cubic feet per minute per person (, ft3/min), given by
where (m3) is the volume of air within the building. Eqs. (8)–(10) are evaluated using the parameter distributions listed in Tabs. S1–S3, and the resulting distribution is then combined with the other methods from this section.
2.5 Fully-developed velocity and pollutant profiles
We employ established formulas to determine fully-developed velocity and pollutant profiles, including the Monin–Obukhov (M-O) similarity theory for natural airflows and approximate solutions of the Navier-Stokes equations for internal airflows, with explanations provided in Section S3.
For external airflows over cities or solar farms, we utilise M-O similarity theory 47, which incorporates the effects of buoyancy and surface roughness under varying thermal conditions. The normal flow rate of pollutants to the surface of the environment is
where represents the average disturbance height and is the roughness length. For , where is the viscous sublayer thickness, we recover the log-law profile. This log law applies to external airflows over smooth or mildly rough surfaces and can be used to calculate pollutant fluxes near vehicles 40.
For internal airflows within HVAC systems, we approximate solutions of the Navier-Stokes equations and boundary conditions 58. The normal flow rate of pollutants to the channel walls is given by
where is the channel diameter. For internal airflows through HVAC filters, we adjust the velocity from the duct flow (with velocity , diameter and cross-channel area ) to the pleat flow (with velocity , diameter , filter sheet area , number ). Flow-through systems (e.g., HVAC filters) are treated as a distinct category of surface-based removal because air is forced through a reactive filter sheet, increasing pollutant–surface encounter frequency relative to diffusive exposure along surfaces. We apply the principle of mass conservation to match average velocities, leading to . The flow rate of pollutants to the pleat channel walls is
The flow rate of pollutants to the fibre sheet is
where is the collection efficiency, related to the permeability and porosity of the filter medium 19. Eqs. (11)–(14) are evaluated using the parameter distributions listed in Tabs. S1–S3, and the resulting distribution is then combined with the other methods from this section.
3 Parameter estimation and uncertainty
To account for variability in atmospheric pollutant transport, we estimated ranges for key parameters across natural, internal and external environments, including characteristic length, surface area, velocity, and total and produced numbers for each environment (Tab. S2–S4). Minimum, mean and maximum values were compiled from published literature and, where available, measurement datasets (e.g., wind speeds in cities 17, Fig. 1E). Where sufficient data existed, we calculated the mean, minima and maxima from the reported values. For parameters with limited data, we adopted a representative value from the literature for the mean and estimated a plausible range, typically spanning approximately one order of magnitude, to reflect environmental variability and reporting uncertainty. Each parameter was represented as a triangular distribution (minimum–mean–maximum) to capture uncertainty, and this distribution was propagated through the atmospheric pollutant surface flux and flow rate equations (1)–(14). This generated a distribution of outcomes, which we illustrate in the Results using box-and-whisker plots, where the central line denotes the median, the box bounds represent the interquartile range (25th to 75th percentiles), and the whiskers indicate the full range of simulated values. Mean values are overlaid for each method to aid comparison.
@@ Results
Natural (e.g., cities, solar farms), internal (e.g., HVAC, combustion, DAC systems) and external environments (e.g., aeroplanes, trains, automobiles) exhibit substantial airflow volumes through their boundary layers, with global medians of m3/s (range: to m3/s) for natural, m3/s ( to m3/s) for internal and m3/s ( to m3/s) for external environments (Fig. 1A). Our analysis of variable parameters including length, surface area, velocity and number of environments, suggests that boundary-layer area and airflow velocity largely determine the flow rate through each environment: cities have the largest total surface area, whereas aircraft have higher velocities (Fig. 1B–G; Tabs. S2–S4). The parameter distributions define the range of flow-rate values and indicate the relative potential of these environments for practical deployment of removal technologies.
@@
The following results describe the atmospheric pollutant fluxes to the surfaces of different environments; these values are determined solely by transport processes and are therefore also independent of the removal technology. Fig. 2A shows normalised atmospheric pollutant fluxes to the surfaces of natural environments (columns 1–2) in cities and solar farms. GHG fluxes exceed other pollutant fluxes due to higher diffusivities and atmospheric concentrations (Fig. 2C). Atmospheric pollutant fluxes to surfaces are influenced by the development length of the boundary layer. Shorter development lengths reduce boundary layer thickness, increase turbulence intensity and enhance the transfer of atmospheric pollutants to surfaces 40. Cities exhibit average surface fluxes of mol/m2s for CO2 (range: to mol/m2s), while solar farms exhibit average surface fluxes of mol/m2s for CO2 ( to mol/m2s) (Fig. 2C, Tab. S2). Although solar farms exhibit higher surface fluxes, the larger total surface area of cities will result in greater pollutant flow rates to their surfaces. Implementing city-scale removal would require integrating removal technologies across millions of square meters of existing surfaces, whereas solar farms could leverage more rapidly expanding infrastructure to integrate the technology as it’s built (Tab. S2).
Using empirical formulae detailed in Methods, average atmospheric pollutant fluxes at the repeating unit scale (e.g., to the surface of individual buildings, m-1) are larger than those at the combined unit scale (e.g., to the surface of entire cities, m-1) due to steeper concentration gradients (Figs. 2A, D, E). Urban turbulence and building spacing will affect pollutant flow rates 6. In Methods, these effects are incorporated via turbulent diffusivity and boundary-layer formulations at city and building scales. The two-dimensional slice shown in Fig. 2D is an idealised illustration, consistent with common approaches for estimating surface fluxes.
@@
Figs. 2A–B highlight the enhanced atmospheric pollutant transport to the surfaces of internal airflows (columns 3–5) in HVAC, combustion, DAC systems (Fig. 2A) and HVAC filters (Fig. 2B). Comparing HVAC, combustion and DAC systems, which show average surface fluxes of , and mol/m2s for CO2 in Fig. 2C, demonstrates that higher airflow velocities enhance atmospheric pollutant transport. The estimated value for DAC is comparable to mol/m2s estimated in McQueen et al. 32. Consistent with our definition of flow-over and flow-through systems (Methods), HVAC filters enhance encounter rates compared with passive duct walls. At the duct wall, the rate at which pollutants move perpendicular to the surface (normal flux) depends on diffusion and is proportional to (e.g., an average of mol/m2s for CO2, Fig. 2C, F), where is the diffusivity of the pollutant, is its concentration and is the height of the channel. In contrast, for HVAC filters, the pollutant flux through the porous fibre sheet is proportional to (e.g., an average of mol/m2s for CO2, Fig. 2C, G), where is the airflow velocity through the fibre sheet. Since and are typically small relative to and , is generally larger than , highlighting the benefits of utilising advective over diffusive transport to improve atmospheric pollutant fluxes to surfaces.
@@
Lastly, we evaluate the normalised atmospheric pollutant fluxes to the surfaces of external airflows over aeroplanes, trains and automobiles (columns 6–8, Fig. 2A). Automobiles demonstrate higher pollutant fluxes to surfaces than trains and aeroplanes (e.g., an average of , , mol/m2s for CO2, Fig. 2C), due to smaller characteristic lengths. However, aeroplanes generate more turbulent transport due to their higher velocities, resulting in a pollutant flux that is approximately 23% of the automobile flux across the entire surface.
@@
The atmospheric surface flow rates presented here are calculated from the fluxes and are independent of the removal method employed. For comparison of total flow to surfaces across applications, we include both flow-over and flow-through categories, even though their underlying transport mechanisms differ. Fig. 3 quantifies the flow rates of atmospheric CO2, CH4, PM2.5 and NO to all global surfaces. These flow rates are evaluated by multiplying the fluxes from Fig. 2 by the relevant surface area for each application ( in Fig. 2D–G) and the number of each application. Natural and internal systems differ in the total pollutant flow to their surfaces, with cities channelling an average of 30 GtCO2/y, three times higher than HVAC systems (10 GtCO2/y) and nearly an order of magnitude above HVAC filters (4 GtCO2/y) (Fig. 3A). While this indicates that city environments offer the largest surface area to achieve gigaton-scale removal targets, realising this potential would require coating surfaces with removal technologies on a massive scale and maintaining their performance over decades. HVAC filters, although receiving lower pollutant flow rates to their surfaces, provide more controlled conditions that may simplify fitting, monitoring and maintenance; moreover, their regular replacement makes accessing the installed base more feasible, making them a promising near-term option. Conversely, external airflows and small-surface-area internal systems (columns 5–8) show lower atmospheric pollutant flow rates to their surface, e.g., an average of 0.1 GtCO2/y for aeroplanes, 0.7 GtCO2/y for trains, GtCO2/y for combustion systems and GtCO2/y for DAC systems, despite external airflows exhibiting higher atmospheric pollutant fluxes to their surfaces in Fig. 2A–C.
To approximate potential deployment rates, we scale atmospheric pollutant transport to yearly produced surfaces rather than total existing surfaces. Because infrastructure is replaced or expands gradually, flow rates to new surfaces decrease: CO2 transport to urban surfaces drops from 30 Gt/y for existing infrastructure to 0.2 Gt/y for new surfaces, while solar farms decline from 0.08 Gt/y to 0.03 Gt/y (Fig. 3; Fig. S2). This contrast arises from differing growth rates: solar capacity is expanding at about 20% per year, compared to approximately 4% annual growth in urban area (Tab. S2). These results indicate that although cities contain the largest absolute removal capacity, solar farms accumulate removal potential more rapidly, underscoring that deployment timescales, not just total capacity, are critical for planning implementation.